# Multi-Domain Evolutionary Optimization of Network Structures

### v1

@@ IV-D Many-network evolutionary optimization
In the evolutionary optimization of community deception, the chromosome is represented by the combination of genes (edges) with a parameter $\rho$ differentiating the edges to add or delete. The population of network $G_{i}\in\mathcal{G}$ can be defined as follows:
$P_{i}=\left[\begin{array}[]{cccccc}e_{1,1}^{i}&e_{1,2}^{i}&\cdots&\cdots&e_{1,\beta_{i}}^{i}&\rho_{1}^{i}\\ e_{2,1}^{i}&e_{2,2}^{i}&\cdots&\cdots&e_{2,\beta_{i}}^{i}&\rho_{2}^{i}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ e_{n,1}^{i}&e_{n,2}^{i}&\cdots&\cdots&e_{n,\beta_{i}}^{i}&\rho_{b}^{i}\end{array}\right],$ (25)
@@
Let $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\ldots,\mathcal{C}^{\prime}_{k^{\prime}}\}$ represent two community partitions. The confusion matrix $m$ is employed to quantify the dissimilarity between these two community partitions. The entry $m_{ij}$ in the matrix denotes the number of common elements between $\mathcal{C}_{i}$ and $\mathcal{C}^{\prime}_{j}$. To facilitate the disengagement of nodes from their initial community affiliations, we can formulate
$F_{1}=-\sum_{i=1}^{|\widetilde{\mathcal{C}}|}\sum_{j=1}^{|\widetilde{\mathcal{C}}^{\prime}|}\frac{|\mathcal{C}_{i}|}{|V|}\left(\frac{m_{ij}}{|\mathcal{C}_{i}|}\log_{2}\frac{m_{ij}}{|\mathcal{C}_{i}|}\right).$ (26)
The objective of Eq. 26 is to foster decentralization within the community structure, albeit it may inadvertently result in numerous tiny communities. Ideally, the nodes from one community should integrate seamlessly into other communities without significantly altering the community sizes. To mitigate this extreme scenario, we introduce a robust alternative function that recalibrates the confusion matrix as follows:
$m^{\prime}_{ij}=\frac{m_{ij}}{|\mathcal{C}^{{}^{\prime}}_{j}|}*\log_{2}|\mathcal{C}^{{}^{\prime}}_{j}|.$ (27)
The first item, ${m_{ij}}/{|\mathcal{C}^{{}^{\prime}}_{j}|}$ refers to the proportion of the nodes of the original community $\mathcal{C}_{i}$ accounting for the new community $\mathcal{C}^{{}^{\prime}}_{j}$. The second term is utilized to prevent situations where a single node forms a community independently. Instead of resorting to computing a global indicator like entropy, we opt for an elegant yet effective approach by constraining its maximum value as follows:
$F_{2}=e^{-max(m^{\prime}_{ij})}.$ (28)
Then, the fitness function $F$ can be defined as follows:
$F=F_{1}*F_{2}.$ (29)
[…]
@@
The transfer solution requires several rounds of reproduction to be integrated into the population of target domain. Combined with the interest of efficiency, we thus take $k=5$ empirically as the interval to determine whether to transfer or not. Let the best fitness values on $G_{i}$ in the current generation and $k$ generations ago be $F^{i}_{n\_0}$ and $F^{i}_{n\_k}$, respectively. Then, we can obtain the fitness difference between the current generation and $k$ generations ago and the fitness difference between $k$ and $2k$ generations ago as follows:
$d_{1}^{i}=\left|F_{n\_0}^{i}-F_{n\_k}^{i}\right|,\quad d_{2}^{i}=\left|F_{n\_k}^{i}-F_{n\_2k}^{i}\right|.$ (31)
@@
$\begin{cases}d_{1,n}^{i}<d_{2,n}^{i},\\ n\geq 2k,\\ n\bmod k=0.\end{cases}$ (32)
The motivation of Eq.32 is that if the evolution speed is slower than the previous, then excellent solutions of other similar domains will be transferred in to facilitate the optimization.
To better use the transfer budget, we propose a self-adaptive mechanism to determine the number of solutions to transfer from $G\in\mathcal{G}^{\mathcal{S}}_{i}$ to $G_{i}$. Each transfer will be evaluated by its contribution to the new population. We will observe the difference between the current elite population $P^{Elite}_{i,n\_0}$ and the elite population before the last transfer $P^{Elite}_{i,n\_l}$, and its intersection with the last transfer ($P^{Transfer}_{i\leftarrow j,n\_l}$) from $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$. That is,
$I_{i\leftarrow j}^{\mathcal{G}}=\frac{|(P^{Elite}_{i,n\_0}\setminus P^{Elite}_{i,n\_l})\bigcap P^{Transfer}_{i\leftarrow j,n\_l}|}{P^{Elite}_{i,n\_0}}.$ (33)
@@
In this manner, we are able to control the ratio of transferred solutions from different networks, given a fixed budget of transfer $\beta^{Transfer}$. The networks that are more similar to the target network should transfer more solutions, i.e.,
$|P^{Transfer}_{i\leftarrow j}|=\beta^{Transfer}*\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (35)
@@
To better use the knowledge of other domains, we design a new mutation operator, which utilizes the elite edge of other networks to replace the edge in its own domain. Given a network $G_{i}\in\mathcal{G}$, the probability of the network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$ being selected to transfer is defined as follows:
$|P^{M}_{i\leftarrow j}|=\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (36)
@@
Note that as different networks have different budgets $\beta$, the repairs to transferred solutions are necessary to ensure compatibility when the sizes of the transferred solution and the solution of the target network are not matched. Given an assisted network $G_{A}$ and a target network $G_{T}$, if $\beta_{A}>\beta_{T}$, then the edges in the transferred solution will be removed randomly until the length of the transferred solution meets the requirement. Otherwise, the edges will be sampled from existing edges as deletion and nonexistent edges as addition to make the transferred solution valid.
In the process of initialization, the chromosomes, denoting edge sets, undergo initialization through a sampling process, where both existing and nonexistent edges are included, with a random ratio. In selection, the roulette wheel strategy is utilized to select the chromosomes to participate in evolution. The chromosome with a larger fitness will be more likely to be selected. The elitism strategy is also adopted to safeguard against the interference of inferior genes, in which the top 10% of chromosomes with superior fitness are retained and directly propagated to the succeeding generation. Note that the operation of crossover and mutation may change the ratio of the added edge and deleted edge in the chromosome (edge set), thus the parameter $\rho$ will be updated during evolution.
@@
Output: The optimized solution for the task on each network
@@
2:Initialize correlation coefficient matrix based on similarity matrix
@@
12:              Update and normalize correlation matrix
@@ V-A Experimental setting
We use eight different real-world networks [53] of various sizes from different domains to validate our proposed MDEO. In the process of evolution, the population size of each network is set to 100. In the reproduction, the probabilities of crossover and mutation are set to 0.5 and 0.1, respectively. In the comparison, the effectiveness of random modification will be investigated and the pioneering work in community deception, DICE [43] will also used as one of baselines. We also optimize the modularity that is the most recognized metric in the studies of community, using the genetic algorithm as in [28] to conceal community information. To demonstrate the effectiveness, the classical evolutionary optimization, referred to single-domain evolutionary optimization (SDEO), will also be included in the comparison.
V-B Experiment on real-world networks
@@
In Table II(b), we compare the efficacy of different community deception algorithms by examining two metrics of NMI and ARI. The lower NMI and ARI are, the better protection the community deception algorithm achieves. We also use Wilcoxon rank sum tests to observe the statistical difference between the compared methods and MDEO, with the significance level $p=0.05$. The symbols ‘+’ and ‘-’ means the compared method is better/worse than MDEO and $\approx$’ means that the compared method has the similar performance with MDEO. The community deception algorithms are tested with three attackers (community detection algorithms) on eight networks of different domains. NMI and ARI basically exhibit the same trend but there are still some exceptions. For example, the NMI value of SDEO and MDEO is higher than DICE while their ARI values are much lower than that of DICE. As observed, our proposed MDEO clearly outperforms other methods in most cases and is better than SDEO, suggested by its lower average NMI and ARI than other compared methods. This result indicates the effectiveness of our method in securing community information. On the other hand, the result is subject to the attacker. For example, the NMI and ARI in Adjnoun when dealing with the attack from FastGreedy, are lower than other networks. Network is also an important variable, as suggested by that NMI and ARI in Adjnoun are higher when dealing with WalkTrap than other networks.
@@
TABLE I: The comparison of the mean and standard deviation of NMI and ARI obtained from different community deception algorithms. The results are averaged from 20 independent runs. Three community detection algorithms are tested, i.e., FastGreedy and WalkTrap.
@@
RAM DICE Mod SDEO MDEO RAM DICE Mod SDEO MDEO
Dolphins 0.75$\pm$0.060(-) 0.74$\pm$0.082(-) 0.76$\pm$0.094(-) 0.51$\pm$0.056(-) 0.43$\pm$0.041 0.70$\pm$0.072(-) 0.69$\pm$0.091(-) 0.71$\pm$0.115(-) 0.31$\pm$0.053(-) 0.24$\pm$0.032
Lesmis 0.79$\pm$0.113(-) 0.75$\pm$0.118(-) 0.80$\pm$0.093(-) 0.55$\pm$0.024(-) 0.47$\pm$0.035 0.72$\pm$0.178(-) 0.65$\pm$0.187(-) 0.76$\pm$0.113(-) 0.31$\pm$0.026(-) 0.25$\pm$0.025
Polbooks 0.85$\pm$0.100(-) 0.89$\pm$0.079(-) 0.92$\pm$0.060(-) 0.54$\pm$0.047($\approx$) 0.52$\pm$0.035 0.84$\pm$0.117(-) 0.89$\pm$0.093(-) 0.92$\pm$0.079(-) 0.45$\pm$0.072(-) 0.41$\pm$0.054
Adjnoun 0.54$\pm$0.118(-) 0.42$\pm$0.094(-) 0.48$\pm$0.118(-) 0.24$\pm$0.036(-) 0.24$\pm$0.026 0.43$\pm$0.159(-) 0.29$\pm$0.119(-) 0.37$\pm$0.154(-) 0.09$\pm$0.018(-) 0.07$\pm$0.022
Erdos 0.68$\pm$0.090(-) 0.67$\pm$0.067(-) 0.70$\pm$0.080(-) 0.50$\pm$0.028(+) 0.49$\pm$0.017 0.58$\pm$0.134(-) 0.58$\pm$0.103(-) 0.61$\pm$0.125(-) 0.28$\pm$0.026(+) 0.27$\pm$0.016
USAir 0.73$\pm$0.114(-) 0.65$\pm$0.126(-) 0.73$\pm$0.111(-) 0.48$\pm$0.057(-) 0.41$\pm$0.045 0.70$\pm$0.168(-) 0.64$\pm$0.163(-) 0.74$\pm$0.139(-) 0.34$\pm$0.055(-) 0.27$\pm$0.050
Netscience 0.95$\pm$0.029(-) 0.93$\pm$0.032(-) 0.96$\pm$0.026(-) 0.83$\pm$0.015(-) 0.80$\pm$0.020 0.90$\pm$0.071(-) 0.85$\pm$0.085(-) 0.91$\pm$0.068(-) 0.57$\pm$0.029(-) 0.53$\pm$0.036
BioCelegans 0.52$\pm$0.058(-) 0.49$\pm$0.042(-) 0.49$\pm$0.067(-) 0.40$\pm$0.036($\approx$) 0.39$\pm$0.029 0.42$\pm$0.084(-) 0.37$\pm$0.073(-) 0.39$\pm$0.010(-) 0.22$\pm$0.033($\approx$) 0.21$\pm$0.024
@@
RAM DICE Mod SDEO MDEO RAM DICE Mod SDEO MDEO
Dolphins 0.67$\pm$0.063(-) 0.64$\pm$0.077(-) 0.67$\pm$0.055(-) 0.57$\pm$0.031(-) 0.54$\pm$0.037 0.51$\pm$0.092(-) 0.48$\pm$0.088(-) 0.51$\pm$0.077(-) 0.30$\pm$0.029(-) 0.27$\pm$0.034
Lesmis 0.90$\pm$0.039(-) 0.86$\pm$0.047(-) 0.90$\pm$0.044(-) 0.81$\pm$0.022(-) 0.74$\pm$0.024 0.80$\pm$0.101(-) 0.72$\pm$0.109(-) 0.77$\pm$0.120(-) 0.55$\pm$0.027(-) 0.48$\pm$0.023
Polbooks 0.92$\pm$0.060(-) 0.88$\pm$0.044(-) 0.90$\pm$0.072(-) 0.64$\pm$0.017($\approx$) 0.63$\pm$0.014 0.93$\pm$0.064(-) 0.88$\pm$0.056(-) 0.89$\pm$0.122(-) 0.49$\pm$0.021($\approx$) 0.48$\pm$0.017
Adjnoun 0.79$\pm$0.035(-) 0.76$\pm$0.051(-) 0.80$\pm$0.042(-) 0.72$\pm$0.022(+) 0.72$\pm$0.018 0.52$\pm$0.111(-) 0.48$\pm$0.150(-) 0.52$\pm$0.126(-) 0.20$\pm$0.035(+) 0.20$\pm$0.022
Erdos 0.90$\pm$0.014(-) 0.87$\pm$0.022(-) 0.88$\pm$0.021(-) 0.85$\pm$0.012(+) 0.85$\pm$0.010 0.70$\pm$0.055(-) 0.62$\pm$0.071(-) 0.63$\pm$0.074(-) 0.47$\pm$0.036($\approx$) 0.45$\pm$0.031
USAir 0.83$\pm$0.045(-) 0.82$\pm$0.066(-) 0.83$\pm$0.041(-) 0.77$\pm$0.014(-) 0.75$\pm$0.009 0.77$\pm$0.070(-) 0.74$\pm$0.139(-) 0.76$\pm$0.103(-) 0.48$\pm$0.044(-) 0.44$\pm$0.036
Netscience 0.94$\pm$0.027(-) 0.92$\pm$0.023(-) 0.93$\pm$0.024(-) 0.86$\pm$0.010(-) 0.85$\pm$0.006 0.82$\pm$0.108(-) 0.75$\pm$0.102(-) 0.80$\pm$0.095(-) 0.53$\pm$0.026(-) 0.51$\pm$0.017
BioCelegans 0.78$\pm$0.029(-) 0.72$\pm$0.028(-) 0.79$\pm$0.030(-) 0.74$\pm$0.011(-) 0.73$\pm$0.009 0.45$\pm$0.060(-) 0.50$\pm$0.063(-) 0.47$\pm$0.081(-) 0.31$\pm$0.017(-) 0.30$\pm$0.017

### v2

@@ IV-D Many-network evolutionary optimization
In evolutionary optimization for community deception, the chromosome is represented by a combination of genes (edges). The population of network $G_{i}\in\mathcal{G}$ can be defined as follows:
$P_{i}=\left[\begin{array}[]{cccccc}e_{1,1}^{i}&e_{1,2}^{i}&\cdots&\cdots&e_{1,\beta_{i}}^{i}&\rho_{1}^{i}\\ e_{2,1}^{i}&e_{2,2}^{i}&\cdots&\cdots&e_{2,\beta_{i}}^{i}&\rho_{2}^{i}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ e_{n,1}^{i}&e_{n,2}^{i}&\cdots&\cdots&e_{n,\beta_{i}}^{i}&\rho_{n}^{i}\end{array}\right],$ (25)
@@
Let $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\ldots,\mathcal{C}^{\prime}_{k^{\prime}}\}$ represent two community partitions. The confusion matrix $m$ is employed to quantify the dissimilarity between these two community partitions. The entry $m_{ij}$ in the matrix denotes the number of common elements between $\mathcal{C}_{i}$ and $\mathcal{C}^{\prime}_{j}$. To facilitate the disengagement of nodes from their initial community affiliations, we can formulate the fitness function as follows:
$F=-\left(\sum_{i=1}^{|\widetilde{\mathcal{C}}|}\sum_{j=1}^{|\widetilde{\mathcal{C}}^{\prime}|}\frac{|\mathcal{C}_{i}|}{|V|}\left(\frac{m_{ij}}{|\mathcal{C}_{i}|}\log_{2}\frac{m_{ij}}{|\mathcal{C}_{i}|}\right)\right)*e^{-\max(m^{\prime}_{ij})},$ (26)
where $m_{ij}$ denotes the number of common nodes between $\mathcal{C}_{i}$ and $\mathcal{C}^{\prime}_{j}$, and $m^{\prime}_{ij}=\frac{m_{ij}}{|\mathcal{C}^{{}^{\prime}}_{j}|}*\log_{2}|\mathcal{C}^{{}^{\prime}}_{j}|$.
For a network $G\in\mathcal{G}$, considering the knowledge transferred from all the rest networks $\mathcal{G}/G$ is undesirable because 1) the optimization efficiency will be compromised with the increase in the networks participating in transfer; 2) the solution from networks that have a low similarity with the target network may bring negative transfer. Therefore, only the solutions from those networks similar to the target network will be transferred to the target population. Let $\mathcal{G}^{\mathcal{S}}_{i}$ be the set of top $\sqrt{|\mathcal{G}|}$ similar networks to network $G_{i}$, which satisfies
@@
The transferred solution requires several rounds of reproduction to be integrated into the population of the target domain. Combined with the interest of efficiency, we thus take $k=5$ empirically as the interval to determine whether to transfer or not. Let the best fitness values on $G_{i}$ in the current generation and $k$ generations ago be $F^{i}_{n_{0}}$ and $F^{i}_{n_{k}}$, respectively. Then, we can obtain the fitness difference between the current generation and $k$ generations ago and the fitness difference between $k$ and $2k$ generations ago as follows:
$d_{1}^{i}=\left|F_{n_{0}}^{i}-F_{n_{k}}^{i}\right|,\quad d_{2}^{i}=\left|F_{n_{k}}^{i}-F_{n_{2k}}^{i}\right|.$ (28)
@@
$\begin{cases}d_{1}^{i}<d_{2}^{i},\\ n\geq 2k,\\ n\bmod k=0.\end{cases}$ (29)
The motivation of Eq. 29 is that if the evolution speed is slower than the previous, then excellent solutions of other similar domains will be transferred to facilitate the optimization.
To better use the transfer budget, we propose a self-adaptive mechanism to determine the number of solutions to transfer from $G\in\mathcal{G}^{\mathcal{S}}_{i}$ to $G_{i}$. Each transfer will be evaluated by its contribution to the new population by observing the difference between the current elite population $P^{Elite}_{i,n_{0}}$ and the elite population before the last transfer $P^{Elite}_{i,n_{l}}$, and its intersection with the last transfer ($T_{i\leftarrow j,n_{l}}$) from $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$. That is,
$I_{i\leftarrow j}^{\mathcal{G}}=\frac{|(P^{Elite}_{i,n_{0}}\setminus P^{Elite}_{i,n_{l}})\bigcap T_{i\leftarrow j,n_{l}}|}{|P^{Elite}_{i,n_{0}}|},$ (30)
where $T_{i\leftarrow j}=\{M_{i\leftarrow j}(s)\mid s\in S_{j}\},$ and $S_{j}$ is a set of selected solutions from network $G_{j}$, and $M_{i\leftarrow j}$ is the function mapping each solution from the domain of $G_{j}$ to that of $G_{i}$.
@@
where $\mathcal{S}^{\mathcal{G}}_{i,j}$ refers to the similarity between $G_{i}$ and $G_{j}$, which can be used to control the ratio of transferred solutions from different networks, given a fixed number of transferred solutions $|T_{i}|$. The networks that are more similar to the target network should transfer more solutions, i.e.,
$|T_{i\leftarrow j}|=|T_{i}|*\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (32)
Elements in the similarity matrix corresponding to non-similar networks are set to zero ($\mathcal{S}^{\mathcal{G}}_{i,j}=0,G_{j}\notin\mathcal{G}^{\mathcal{S}}_{i}$) prior to optimization, whereas elements representing network similarity are normalized during each update.
Let the population of network $G_{i}$ be $P_{i}$ and knowledge transfer occurs from similar networks $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$, and the transferred high-quality solution $T_{i\leftarrow j}$ from a network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$ is incorporated into the population to assist the optimization process. Therefore, the updated population is given by:
$P_{i}=\begin{cases}P_{i}\cup\bigcup_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}T_{i\leftarrow j},&\text{if knowledge transfer occurs,}\\ P_{i},&\text{otherwise.}\end{cases}$} (33)
@@
To better utilize the knowledge of other domains, we introduce a new mutation mechanism that leverages elite edges from other networks to replace edges in the current domain. The probability of selecting a network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$ for knowledge transfer is defined as follows:
$\mathbb{P}^{M}_{i\leftarrow j}=\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (34)
Given a network $\mathcal{G}_{i}$, the deletion and addition candidates are denoted by $E_{i}^{M,D}$ and $E_{i}^{M,A}$, respectively. When knowledge transfer is applied, these candidates are derived from the transferred solutions of assisted networks $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$, which are organized into addition ($T_{i\leftarrow j}^{A}$) and deletion ($T_{i\leftarrow j}^{D}$) sets. The mutation candidate sets are then updated accordingly as follows:
$E_{i}^{M,A}=\bigcup_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}T_{i\leftarrow j}^{D},\quad E_{i}^{M,D}=\bigcup_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}T_{i\leftarrow j}^{A},$} (35)
where $T_{i\leftarrow j}^{A}$ represents the set of edges for addition, and $T_{i\leftarrow j}^{D}$ represents the set of edges for deletion from the transferred solution of network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$.
As for the selection in our work, each individual has an equal probability of being selected, promoting diversity and helping prevent premature convergence on suboptimal solutions. Additionally, we adopt an elitist mechanism to ensure that the highest-quality solutions are preserved.
@@
As different networks have different budgets $\beta$, the repairs to transferred solutions are necessary to ensure compatibility when the sizes of the transferred solution and the solution of the target network are not matched. Given an assisted network $G_{A}$ and a target network $G_{T}$, if $\beta_{A}>\beta_{T}$, then the edges in the transferred solution will be removed randomly until the length of the transferred solution meets the requirement. Otherwise, the edges will be sampled from existing edges as deletion and nonexistent edges as addition to make the transferred solution valid. The pseudocode of the many-network evolutionary optimization can be found in Algorithm 3.
@@
Output: The optimal solutions for the task across each network
@@
2:Initialize similarity matrix
@@
12:     Update and normalize similarity matrix
@@
IV-E Complexity analysis
[…]
If the networks are optimized independently, the time complexity is $O(|\mathcal{G}|\cdot|P|\cdot N_{t})$, where $N_{t}$ denotes the number of evolutionary rounds, and $|P|$ is the population size in each network. In the MDEO framework, the solutions from $\log|\mathcal{G}|$ networks are transferred to assist others, but the number of solutions transferred, denoted by $|T|$, is fixed, and all assisted networks share this budget. Consequently, the worst-case complexity for MDEO could be $O(|\mathcal{G}|\cdot|P\cup T|\cdot N_{t})$. However, transfers only occur at least every five generations and additionally must satisfy the requirements of Eq. 29. Therefore, the practical complexity will be lower than this worst-case scenario. Since $|T|$ is a constant and significantly smaller than $|P|$, the complexity of MDEO will only be marginally higher than that of separately optimized networks, i.e., single-domain evolutionary optimization (SDEO), thereby justifying the practical efficiency of MDEO.
@@ V-A Experimental setting
To obtain graph embeddings, we adopt a two-layer GCN as the encoder where the degree centrality, closeness centrality and community information are collected as node features. In the process of evolution, the population size of each network is set to 100, and the probabilities of crossover and mutation are set to 0.5 and 0.1, respectively. The number of transferred solutions is 30. In our study, we use two classical community detection algorithms, FastGreedy [62] and WalkTrap [63], as attackers to assess whether the community structure has been effectively obfuscated through community deception methods. The results are averaged over 20 independent runs and the implementation data is available online [^1] https://networkrepository.com/.
V-B Benchmark
The mainstream approaches to community deception include heuristic methods, GNN-based method, and evolutionary optimization techniques. Accordingly, some of each have been chosen as benchmarks, specifically: RAM: A heuristic approach that randomly rewires all possible edges in the network [51]; DICE: A heuristic approach that randomly deletes intra-community edges and adds inter-community edges [51]; Mod: This method uses the genetic algorithm to optimize modularity to modify the network edges to conceal community information [27]; CD-ATTACK: This method alters links by using constrained adversarial graphs generator based on graphs neural networks (GNNs) [64]; CoeCo: This approach employs a divide-and-conquer strategy for scalable evolutionary optimization. It optimizes both the subgraphs and the original graph cooperatively to identify the optimal edge set for community obfuscation [30]; SDEO: The classical evolutionary optimization without knowledge transfer, where each network is optimized separately, providing a baseline for comparison against our proposed MDEO.
V-C Experiment on real-world networks
@@
To examine the efficacy of different community deception algorithms, we adopt two metrics of normalized mutual information (NMI) [65] and adjusted rand index (ARI) [66] to observe the change in community structure, shown in Table IV(b). The lower the NMI and ARI are, the better protection the community deception algorithm achieves. The statistical disparities between MDEO and other methodologies were assessed using the Wilcoxon rank sum test, setting the significance level at $p=0.05$. Symbols ‘+’ and ‘-’ are used to show whether the alternative methods outperform or underperform compared to MDEO, respectively, and $\approx$’ signals performance equivalence. The community deception algorithms are tested with two attackers (community detection algorithms) on eight networks of different domains. As observed, our proposed MDEO consistently outperforms other methods, including heuristic approaches (RANDOM, DICE), metaheuristic algorithms (MoD, CoeCo), and the GNN-based method (CD-ATTACK). It also demonstrates superior performance compared to SDEO, as indicated by its lower average NMI and ARI scores across the evaluated networks. This result indicates the effectiveness of our method in securing community information. On the other hand, the result is subject to the attacker. For example, the NMI and ARI in Adjnoun when dealing with the attack from FastGreedy, are lower than other networks. Network is also an important variable, as suggested by that the NMI and ARI in Adjnoun are higher when dealing with WalkTrap than other networks.
@@
TABLE III: The comparison of the mean and standard deviation of NMI and ARI obtained from different community deception algorithms on eight tested networks. FastGreedy and WalkTrap are used to examine the difference in community structure before and after modification, respectively.
@@
RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO
[…]
@@
RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO
[…]